OpenStudy (annoymous697):

Which is the area of a circle with a circumference of 37.68 units? (use 3.14 for )

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (annoymous697):

are u suppose to divide the 37.68 befor you can solve ???

OpenStudy (skullpatrol):

$$\Huge A= \pi r^2$$ $$\Huge C=2 \pi r$$

OpenStudy (annoymous697):

ok thank you

OpenStudy (skullpatrol):